# Simulation as Supervision: Mechanistic Pretraining for Scientific Discovery

### v2

@@ 1 Abstract
Scientific modeling faces a tradeoff: mechanistic models provide scientific grounding but struggle with real-world complexity, while machine learning models achieve strong predictive performance but require large labeled datasets and are not interpretable. We introduce Simulation-Grounded Neural Networks (SGNNs), which use mechanistic simulations as training data for neural networks. SGNNs are pretrained on synthetic corpora spanning diverse model structures, parameter regimes, stochasticity, and observational artifacts. Simulation-grounded learning has been applied in multiple domains (e.g., surrogate models in physics, forecasting in epidemiology). We provide a unified framework for simulation-grounded learning and evaluated SGNNs across scientific disciplines and modeling tasks. We found that SGNNs were successful across domains: for prediction tasks, they nearly tripled COVID-19 forecasting skill versus CDC baselines, reduced chemical yield prediction error by one-third, and maintained accuracy in ecological forecasting where task-specific models failed. For inference tasks, SGNNs also accurately classified the source of information spread in simulated social networks and enabled supervised learning for unobservable targets, such as estimating COVID-19 transmissibility more accurately than traditional methods even in early outbreaks. Finally, SGNNs enable back-to-simulation attribution, a form of mechanistic interpretability. Back-to-simulation attribution matches real-world observations to the training simulations the model considers most similar, identifying which mechanistic processes the model believes best explain the observed data. By providing a unified framework for simulation-grounded learning, we establish when and how mechanistic simulations can serve as effective training data for robust, interpretable scientific inference.
@@ 2 Introduction
Recent work has sought to bridge this divide by incorporating mechanistic knowledge into machine learning. These approaches can be understood through the three core components of machine learning problems: the data, the model architecture, and the optimization procedure. Physics-Informed Neural Networks (PINNs) incorporate governing equations into the loss function, constraining optimization to satisfy known physical laws [1, 2]. Neural ODEs embed continuous-time dynamics directly into the architecture, parameterizing derivative functions that are solved during forward passes [3]. Universal Differential Equations (UDEs) combine both approaches, using neural networks to learn unknown terms in otherwise known equations while training on trajectory data [4]. These methods perform well when equations are known and all relevant variables are observable, but can introduce bias when mechanistic models are incomplete or when systems undergo regime shifts.
@@
We introduce Simulation-Grounded Neural Networks (SGNNs) as a general framework for training neural networks on mechanistic simulations. While simulation-grounded learning has been applied in isolated contexts, we provide a unified treatment and evaluation across domains. SGNNs are pretrained on synthetic corpora spanning diverse model structures, parameter regimes, stochasticity, and realistic observational artifacts. This approach treats mechanistic models as generators of labeled training data, allowing neural networks to internalize scientific structure without requiring it to be incorporated into architectures or loss functions. This gives a flexible framework compatible with any architecture or optimization method.
@@
These results establish SGNNs as a general framework for scientific machine learning, demonstrating when and how mechanistic simulations can serve as effective training data for robust, interpretable inference across disciplines. [^1] Code and data available at https://github.com/carsondudley1/SGNNs.
@@ 5 Discussion
Simulation-grounded learning bridges mechanistic and empirical modeling by embedding scientific structure through examples rather than architectural constraints or loss functions. This approach is compatible with any architecture and optimization method, supporting transfer learning, fine-tuning, and straightforward integration with existing deep learning infrastructure. Because simulations span diverse model structures and parameter regimes, SGNNs can be robust to structural mismatch between training simulations and real-world systems.
Training on simulated data provides ground-truth labels for quantities that are unobservable in real settings, like transmission rates or carrying capacities that lack direct measurements. This enables approaches like simulation-based inference (SBI), which use supervised learning for parameter inference. Our results show that various approaches are possible within this framework: from full Bayesian posterior estimation, established with SBI, to simpler methods like quantile regression, which can achieve accurate parameter inference with reduced computational complexity.
Back-to-simulation attribution links real-world inputs to training simulations the model considers similar, revealing which mechanistic regimes the model associates with observed patterns. Because each retrieved simulation has known ground-truth parameters and generative processes, users can inspect what the model believes is happening mechanistically. For example, if a model retrieves simulations with high asymptomatic transmission when shown an outbreak with few reported cases, this suggests the model interprets the pattern as indicating substantial hidden spread. Unlike feature attribution methods that identify which input variables influenced a prediction, back-to-simulation attribution explains which scientific processes the model believes are active. This provides actionable insight. If retrieved simulations consistently show specific parameter regimes or structural features, practitioners can validate these mechanistic hypotheses against domain knowledge or use them to guide intervention strategies.
Our results suggest that simulation-grounded learning can capture generalizable representations of scientific structure. Models trained only on human-to-human transmission accurately predicted vector-borne dengue dynamics, indicating that neural networks can internalize abstract principles that transfer across mechanistically distinct settings. This points toward the potential for simulation-pretrained foundation models in science, though substantial work remains to understand the scope and limits of such generalization.
@@
SGNNs demonstrate that mechanistic simulations can serve as effective training data across diverse scientific domains, providing a framework for when and how this approach succeeds. By positioning simulation-grounded learning within the broader landscape of hybrid modeling approaches, we establish design principles for combining scientific theory with deep learning flexibility.

### v3

@@ 1 Abstract
Scientific modeling faces a tradeoff between the interpretability of mechanistic theory and the predictive power of machine learning. While hybrid approaches like Physics-Informed Neural Networks (PINNs) embed domain knowledge as functional constraints, they can be brittle under model misspecification. We introduce Simulation-Grounded Neural Networks (SGNNs), a framework that instead embeds domain knowledge into the training data to establish a structural prior.
By pretraining on synthetic corpora spanning diverse model structures and observational artifacts, SGNNs learn the broad patterns of physical possibility. This allows the model to internalize the underlying dynamics of a system without being forced to satisfy a single, potentially incorrect equation.
We evaluated SGNNs across scientific disciplines and found that this approach confers significant robustness. In prediction tasks, SGNNs nearly tripled COVID-19 forecasting skill versus CDC baselines. In tests on dengue outbreaks, SGNNs outperformed physics-constrained models even when both were restricted to incorrect human-to-human transmission equations, demonstrating that SGNNs are potentially more robust to model misspecification. For inference, SGNNs extend the logic of simulation-based inference to enable supervised learning for unobservable targets, estimating early COVID-19 transmissibility more accurately than traditional methods. Finally, SGNNs enable back-to-simulation attribution, a form of mechanistic interpretability that maps real-world data back to the simulated manifold to identify underlying processes. By unifying these disparate simulation-based techniques into a single framework, we demonstrate that mechanistic simulations can serve as effective training data for robust scientific inference that generalizes beyond the limitations of fixed functional forms.
@@ 2 Introduction
Recent work has sought to bridge this divide by incorporating mechanistic knowledge into machine learning. These approaches can be understood through the three core components of machine learning problems: the data, the model architecture, and the optimization procedure. Physics-Informed Neural Networks (PINNs) incorporate governing equations into the loss function, constraining optimization to satisfy known physical laws [1, 2]. Neural ODEs embed continuous-time dynamics directly into the architecture, parameterizing derivative functions that are solved during forward passes [3]. Universal Differential Equations (UDEs) combine both approaches, using neural networks to learn unknown terms in otherwise known equations while training on trajectory data [4]. These methods act as functional constraints: they perform well when equations are known and all relevant variables are observable, but can introduce bias or cause the learning process to fail when mechanistic models are incomplete or structurally misspecified.
@@
We introduce Simulation-Grounded Neural Networks (SGNNs) as a general framework for training neural networks on mechanistic simulations. While simulation-grounded learning has been applied in isolated contexts, we provide a unified treatment and evaluation across domains. Unlike approaches that enforce pointwise constraints, SGNNs embed domain knowledge into the training data to establish a structural prior. By pretraining on synthetic corpora spanning diverse model structures, parameter regimes, and realistic observational artifacts, the model internalizes the global geometry of physically plausible dynamics. This allows the neural network to inherit scientific structure from the underlying math without being explicitly constrained by a single, potentially incorrect governing equation. This resulting framework is architecture-agnostic and maintains flexibility across diverse scientific domains.
@@
These results characterize SGNNs as a general framework for scientific machine learning, demonstrating when and how mechanistic simulations can serve as effective training data for robust, interpretable inference across disciplines. [^1] Code and data available at https://github.com/carsondudley1/SGNNs.
@@ 5 Discussion
Central to this framework is the shift from functional constraints to structural priors. While existing hybrid modeling approaches (e.g., PINNs or UDEs) enforce domain knowledge by explicitly constraining a model’s derivative or loss function, simulation-grounded learning embeds scientific structure through examples. This allows the neural network to internalize the global manifold of physically plausible dynamics without being tied to a single, potentially incorrect functional form. Because simulations span diverse model structures and parameter regimes, SGNNs develop a form of ensemble robustness, generalizing effectively even when real-world dynamics deviate from any single training simulator.
Training on simulated data also provides ground-truth labels for quantities that are unobservable in real settings, like transmission rates or carrying capacities that lack direct measurements. This enables approaches like simulation-based inference (SBI), which use supervised learning for parameter inference. Our results show that various approaches are possible within this framework: from full Bayesian posterior estimation, established with SBI, to simpler methods like quantile regression, which can achieve accurate parameter inference with reduced computational complexity.
Back-to-simulation attribution further enhances the utility of this framework by linking real-world inputs to training simulations the model considers similar. Unlike post-hoc feature attribution methods, this reveals which mechanistic regimes the model associates with observed patterns. For example, if a model retrieves simulations with high asymptomatic transmission when shown an outbreak with few reported cases, it suggests the model interprets the pattern as indicating substantial hidden spread. This provides actionable, process-level insight, allowing practitioners to validate mechanistic hypotheses against domain knowledge or use them to guide intervention strategies.
Our results suggest that simulation-grounded learning can capture generalizable representations of scientific structure that extend beyond the training distribution. The success of this paradigm has been further validated through the development of Mantis [36], a simulation-grounded foundation model for infectious disease forecasting. By scaling the SGNN framework to a massive synthetic corpus, Mantis achieves performance competitive with state-of-the-art models across diseases and geographies entirely out-of-the-box, without requiring real-world fine-tuning. This underscores the potential for simulation-pretrained models to serve as inference engines for science, though substantial work remains to understand the theoretical bounds of such generalization.
@@
SGNNs demonstrate that mechanistic simulations can serve as effective training data across diverse scientific domains. By positioning simulation-grounded learning as a robust alternative to constraint-based hybrid methods, we establish a design principle for combining scientific theory with deep learning flexibility: use mathematical simulators to define a structural prior over the system dynamics, and let the data-driven model learn to perform inference within those bounds.